Status Report

ISS On-Orbit Status 8 Aug 2002

By SpaceRef Editor
August 8, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Crew preparations continued for the EVA-7 and EVA-8 spacewalks on 8/16 and 8/23.  [CDR Korzun and FE-2 Treschev held a TV conference via Ku- and S-band with TsUP (MCC-Moscow) personnel to downlink and review the video shot yesterday of their EVA equipment preparations.  With FE-1 Whitson, they assembled a combined equipment bundle around the fully loaded EVA tool carrier (KPU), including the new Kromka-2 tablet and a semi-rigid airtight container (PGK) for the Kromka-1 experiment to be retrieved from the Service Module (SM) hull.  Since the ham radio antennas WA1 and WA2 will be taken outside and installed during the second spacewalk (EVA-8), they were not included yet.  Later, Korzun and Whitson prepared Orlan component spares as well as auxiliary and individual crewmember equipment.]

During her (successful) switch nut verification, battery installation and checkout activity on the three EVA PGTs (pistol grip tools) yesterday, Peggy Whitson reported a loose bayonet fixture on one unit (#1006).  If the other two PGTs (#1004 & 1008) check out OK, she can use either one of them for EVA-7.  Meanwhile, MCC-H is working on a plan for the repair of PGT #1006, to be task listed later.

MCC-H is also developing a planfor battery discharge troubleshooting in a few days.  [Current recommendation is to partially discharge an EMU (extravehicular mobility unit, US spacesuit backpack) battery in the EMU using the EMU’s internal fan for about an hour.  This would reverse the passivation
(oxalate crystal growth between cells of the battery) and put the battery in a better configuration for discharging in the BSA (battery stowage assembly) using the BCA (battery charger assembly).  The fallback procedure would be to fully discharge the EMU batteries in the EMU using the EMU fan.]

To support transfer and installation of the DPP micrometeoroid/orbital debris shield panels by Valery and Peggy during EVA-7, TsUP is preparing to uplink a multimedia training aid with photos and computer models illustrating these activities.

The crew completed another session of robotics operations, maneuvering the SSRMS (space station remote manipulator system) to the viewing position for EVA-7 and providing a training session for FE-2 Treschev on
controlling its cameras.  [After FE-1 Whitson re-installed the UOP (utility operations panel) bypass cable to the Lab RWS DCP (robotics workstation/display control panel), the ground powered the MSS (mobile servicing system) up, with the MBS (mobile base system) redundant string and SSRMS prime string on Operational.  VDS (video distribution system) was configured from the ground, and the crew then had about an hour to perform the robotics ops.  Afterwards, the MSS was powered down, with Lab RWS off, and SSRMS and MBS on keep-alive power on both strings].

FE-2 Sergei Treschev completed the periodic replacement of the PS1 and PS2 dust collector filters in the FGB.

Sergei also serviced the SM’s harmful impurities/micropurification unit (BMP), starting the bake-out (regeneration) cycle on absorbent bed #1, while channel #2 continues in Purify mode..  Each regeneration cycle takes about 24 hours.

CDR Korzun and FE-1 Whitson underwent their pre-EVA PFE (periodic fitness evaluation), during which they performed prescribed exercise work while the heart rate is logged with the HRM (heart rate monitor), and using equipment from their ambulatory medical and advanced life support packs.  [Next Saturday, 8/10, Valery and Peggy are also scheduled for the Russian MedOps MO-5 assessment (cardiovascular evaluation during graded exercise on the VELO), a standard pre-EVA medical test.]

The fourth SUBSA sample (SUBSA-04) run was completed today, with its thermal profile finished at 5:00 am EDT. 
[SUBSA-04 is a tellurium-doped non-baffle sample.  The MSG (Microgravity Science Glovebox) rack was powered down later in the day after Peggy Whitson removed the specimen from the furnace inside the MSG, and the ground downlinked the SUBSA data from the micro drive.] 

Sample collection ended for FE-2 Treschev on the current NASA/JSC Renal (Kidney) Stone research session and continues, for one more day, for FE-1 Whitson, with metabolic log entries for both subjects.

Elektron update:  CDR Korzun’s work schedule today had 70 minutes blocked out, in several segments, for
recovering the Elektron oxygen generator.  [As reported before, the electrolysis machine is not working properly due to a malfunctioning upper limit switch for the buffer tank, which prevents the solenoid valve from closing when the tank is full.  Today’s recovery attempts by Korzun had him shut down the Elektron and then go through a number of troubleshooting steps, operating valves and using laptop resets and commands, to restore the “Buffer Tank is Full” flag for automatic closure of the solenoid valve.]

TVIS treadmill update:
  A safety analysis is currently in progress to determine how much the treadmill belt would deflect if the chassis truss, wthout the damaged roller, were to fail during exercise.  A Go for resuming crew exercise is expected by tomorrow morning, but with constraints imposed for operation and other activities required for TVIS use (including regular status checks and inspections with photos, etc).

Troubleshooting is continuing on the three BCRs (barcode readers) on board.  [Ground specialists are unable to reproduce the problem reported by the crew in its entirety.  A task list item has been added to tomorrow’s schedulefor Peggy to videotape what she is seeing and then downlink the video for inspection.]

Daily routine servicing tasks completed by CDR Korzun today were SOSh life support systems maintenance and IMS (inventory management system) “delta” file preparation, while FE-1 Whitson conducted the autonomous payloads status checks.

All crewmembers performed their physical exercise program.

The US CEO (crew earth observation) program had the following targets:  Industrialized Southeastern Africa (a cold, dry air mass has settled over the South African interior valleys and this favors smog formation there.  Of interest as ISS tracked northeastward along the coast: looking well inland beyond the coastal ranges to detect industrial smog there), Angolan Biomass Burning (this was a fine pass for oblique views of northern Namibia and most of Angola to the left of track.  Burning season is past its peak now, but both active fires and fire scars are available targets in this region), Congo-Zimbabwe Biomass Burning (the best views of western and northern Zimbabwe were to the right of track this pass.  The dry season will end soon, so the crew was asked to look for fires, fire scars, and vegetation patterns), High Central Andean Glaciers (of interest: taking advantage of the fair weather and good light over this target to acquire context views of the region for use later.  Also: looking to the left of track [NW] along the spine of the Andes to locate the small ice fields and glaciers
, Eastern Mediterranean Dust (a complex storm system over central and western Europe is producing strong westerly winds over northern Africa.  Dust from Algeria, Tunisia, and Libya, visible in satellite imagery and is moving into the eastern Med basin.  On this pass, as the ISS tracked towards Greece, the crew was to look right of track to document this dust plume in oblique and limb views).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:05 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered Off (see above).  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber is Off (leak).  BMP Harmful Impurities unit: Absorbent bed #1 in Regeneration mode, bed #2 in Purify mode.

SM Working Compartment: 
Pressure (mmHg) — 758, temperature (deg C) — 26.3, ppO2 (mmHg) — 156.1, ppCO2 (mmHg) — 2.6.
SM Transfer Compartment:  Pressure (mmHg) — 761, temperature (deg C) — 20.7.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 22.3.
Node:  Pressure (mmHg) — 751.52, temperature (deg C) — 25.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 753.40, temperature (deg C) — 24.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 753.40, temperature (deg C) — 28.1; shell heater temp (deg C) — 27.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 24.5
PMA-2:  Shell heater temp (deg C) — 14.6

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [ a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B also in Autotrack mode.
SM batteries: Battery #1 is off-line, battery #8 is cycling; all other batteries (6) are in “Partial charge” mode.
FGB batteries:  Battery #5 is off-line, battery #4 is cycling; all other batteries (4) are in “Partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs online.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
MBS:  Keep Alive power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:23 am EDT [ epoch]):
Mean altitude — 395.4 km
Apogee — 407.6 km
Perigee — 383.2 km
Period — 92.5 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.001801
Orbits per 24-hr. day — 15.57
Altitude decrease — 100 m (mean) in last 24 hours
Solar Beta Angle — 19.2 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 21218
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane “sun-fixed” [yaw: ~180 deg, pitch: -5.7 deg., roll: 0 deg]).  Next maneuver to LVLH (“earth-oriented”) on Friday, 8/9.

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.